Minutes of the 15th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson

Minutes taker Mark Dennehy

Present Steven Watterson (Captain), Benjamin Jones (Treasurer), Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer), Conor McDermottroe (Public Relations Officer), Peter Byrne, Deirdre Ninaber (Ladies’ Coordinator)

Absent(with apologies)

Absent

Date 27 Nov 2001 1845 – 2100

Distribution DURC Committee

Overview of topics

1 Secretarys Report 1

2 Treasurers Report 2

3 CRO Report 2

4 A&E Report 3

5 DURC 10m Open Review 4

1 Secretarys Report

5.1 Notes for next year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

6 RO Dinner 5

7 Turkey Shoot 5

8 National Squad 5

9 Christmas lists 5

10 Capital Grant 6

11 Any Other Business 6 11.1 Jackets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 11.2 Pony Club coaching . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 11.3 Student Newspapers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 11.4 Website . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 11.5 AGM & Constitutional reform . . . . . . . . . . . . . . . . . . . . . . 7 11.6 Shooting nights . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 11.7 Code of Conduct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

1 Secretarys Report

Mark reported that he had been contacted by Nick West in regard to post delivered to 69 Regents House, apparently an old Rifle Club address. Mark had never heard of this, or seen where it was. Ben said most RRPC mail made it to DUCAC so some internal college routing must take place. Mark agreed to look into this. Task: Look into the Regents House address (Mark)

Mark reported that the NSRA/Eley postal as well as the NSRA Summer season postal competition entry forms had arrived. The deadline for the NSRA/Eley entrys was Jan 16. It was decided to set minimum average scores for shooters wising to enter, 85/100 for prone 25yd, and 70 for 10m Air rifle. The Summer season postals were discussed. It was decided not to enter large numbers of competitions due to the lack of shooters during summer due to holidays. Mark pointed out that 10m shooting falls off out of season and that the summer postal was an ideal opportunity to at least slow this. Mark will recruit shooters for these competitions. Task: Recruit shooters for the postal competitions (Mark)

2 Treasurers Report

Ben reported that we now have £400 in the bank after bills are accounted for, though he suspected that was high and that we had some bill somewhere under the radar. Squad

3 CRO Report

training money has been applied for, but we need bills to claim it. Mark will get Pat O’Brien to send us a bill from the NTSA. John thought that the RO Training Course should be paid for from coaching, Ben said no, it should come from fees because we get very little money for coaching and we can’t claim National Squad training from anything but coaching.

Ben reported that Cathy has said that there is absolutely no way DUCAC will pay for the colours to be held in WTSC. Steven asked what their justification was. Ben said that DUCAC say that every other club hosts colours for free. WTSC have to charge for the new range. UCDRC charged us range hire last year. The question of whether or not to use WTSC’s new range came up and was quickly dismissed as being too expensive. John wondered what the state of the old WTSC range was and Mark reported it had been refurbished and was still in daily use. Ben asked for a decision - UCDRC or WTSC. John was for WTSC for the variety, lack of a home team advantage to anyone and the social aspect of the event. Alan pointed out that we wanted to run the triad competition the next day. Ben wondered if we could stay overnight in WTSC. It was thought it wouldn’t be a problem. Steven wondered if UCD would be better as it was easier to get home from. Alan pointed out that that was a disadvantage from our point of view. Logistics were discussed - people will travel by bus, Mark will drive the equipment out in his car, as he does not expect to stay overnight, claiming old age as a reason. A vote was take on holding the colours in WTSC, all were in favour.

Ben asked if more pellets were needed, John replied that they were.

3 CRO Report

Alan began with the recent problem with the range keys going missing temporarily. John gave the full story, and pointed out that having the safe open was completely unacceptable. Steven pointed out that DUCAC is contacted by college security when this happens and lots of running about ensues. John said that all ROs must supply telephone numbers - home, work, mobile and emergency contact numbers. These should be kept private for the club officers and CRO only. Alan pointed out that David Gallagher (who had the keys) has no mobile at present. He wondered if all ROs should have a mobile as a mandatory requirement. It was decided that this wasn’t workable. Mark brought up Alan’s idea of a duplicate keyset. This was dismissed as being a bad security risk. Steven wondered if Security should have a second set of keys. John pointed out that this would get Security unnecessarily involved before we had the opportunity to recover the keys. Peter wondered if a keyhook at the safe would be a better idea as the current keyhook is in an akward position. John thought that if our current one doesn’t work, a new one wouldn’t either. Steven agreed that the current location was daft. Peter thought that a keyhook in a convienent location would be more likely to be used. Peter suggested a more obvious location, but Ben pointed out that this was a security risk. Deirdre suggested inside the safe door as a location.

Alan wondered why noone checked the logbook in Front Arch. John agreed, and said that Ray had been irresponsible in that he had called John, who was not responsible,

4 A&E Report

and effectively said that he (Ray) had to go home and that John was not in charge of the situation. Alan agreed that this was not on. John was unhappy at Ray’s not checking the logbook for his own expediency. Alan pointed out that it would be hard to check the logbook without security getting upset. Mark pointed out that Ray did call, but didn’t check the book. Peter at this point said that if the committee felt that disciplinary action was called for, we should get Ray in the discussion. Alan said that the committee had to decide if that was called for first. Steven said that there were two courses of action under consideration. Firstly was a review of the keys procedures, and second was disciplinary action against David and Ray. John said that the keys procedure must be reviewed, and Range Officer retraining is required in procedures and form-filling. Deirdre agreed, saying that there was too much covered in the original RO meeting, and another was needed. Steven said that one should be arranged. Ben asked John what he thought Ray should have done. John said he should have taken responsibility as senior RO on the scene and that calling John (who is not CRO) out of bed so Ray could go home was not the correct action. Alan said we needed coherent procedures and a set of contingency plans. Steven said that in principle, this was a serious breach of security, but in practise it wasn’t as serious as first seen. Alan pointed out that until the keys are located, we can’t assume they haven’t been stolen. Mark pointed out that even the appearance of a security risk was a very serious threat to the club. Steven asked for a course of action. Deirdre reitereated her opinion that another RO meeting was needed. Peter said that the importance of procedures needed to be stressed there. Mark said that this loss of keys has happened far too often allready. Conor suggested stressing procedures. Ben said that that wouldn’t work, as it hasn’t in the past. Alan suggested a hardware solution, with a keyswitch for the target lights that forces the keys to be in one position when shooting so their removal is immediately obvious. Mark and Alan then had a technical discussion on how to do this. It was accepted by all after this. Ben wanted to know how much it would cost. Alan is to source and price hardware. John said this should go in the Risk Assessment Document. Task: Source and Price hardware for keyswitch (Alan)

4 A&E Report

John reported on the ammo count discrepencies. ROs need to be more careful with paperwork. 48 T22 rounds are still unaccounted for. Steven and Conor both thought it might be their fault. Mark pointed out that this was a safety risk. Peter wondered if it could just be an accounting mistake. John said we have no way of knowing as there’s no record of the ammo leaving Front Arch or being used on the Range. John said it’s most likely been used in competition and there’s been no crime, but all rounds can be traced to us if used in a crime. Mark pointed out that that sounds funny but definitely is not. Ben pointed out that the number on the box is not always the number of rounds in the box. John said this practise of writing the number of rounds on the outside of the box must cease immediately. All ROs must physically count all rounds. Physical counts must be done whenever time permits. In case a discrepency is found it

5 DURC 10m Open Review

must be immediately reported to John. The ammo procedures must be reformalised as we have been getting far too informal lately, which is an unacceptable state of affairs. Security must be tightened up. ROs must be brought together for a discussion on this and retraining.

Ben wondered when this meeting was scheduled for. Alan set the date at Tuesday Janurary 8th on the range at 1830. Task: Announce RO Meeting (Alan)

John reported we have plenty of targets and .22 ammo, we need some pellets and Rifles S and V are due back from Dave Cooney in a few days.

Steven asked about the behavioual problems with the other rifles (the 603’s trigger and T’s action). Alan said the 603 trigger problem (premature release of the shot) was caused by the trigger being set too lightly and the action being handled too roughly. Peter and John disagreed with this evaluation. Steven thought a servicing was a good idea. Mark asked when the rifles were last serviced, Alan said it had been over two years. Conor pointed out that .22 rifle #4 also misfires. Mark said that he had found the problem with the bolt some months previously and had cleaned the bolt with solvent to keep it working for a few days as a stopgap measure. We were to have given it to Dave Cooney for servicing within the week - it had still not been sent as of this meeting. John has examined the bolts of most of the rifles and most need servicing. #4’s bolt’s firing pin needs sharpening.

John will adjust the 603 trigger to try to solve the problem - if this doesn’t work, we’ll send it to Dave Cooney. The others shouldn’t be all sent off at once. Alan said not to play with the trigger but to get it properly fixed. Ben pointed out that the 603 was now a general-purpose club rifle and shouldn’t be set with a light trigger. Steven said that club rifles shouldn’t be tweaked by experienced shooters to the point where they become a hazard for novice shooters. Task: Adjust 603s trigger and report (John)

5 DURC 10m Open Review

Ben reported a £1500 profit, quickly corrected to a £55 profit. Mark has been paid £20 for his cheesecakes, £20 went on first round drinks, £13 was set aside for engraving the shield and £37 for the NTSA levy.

Mark and Alan settled their entry fees.

Steven wondered about paying UCD a range fee. John wondered about paying Cahal personally as he’d made considerable effort to get the range open for us. This was discussed and it was agreed to pay UCD £20 towards consumables used and to give Cahal a personal present - a nice bottle of wine was chosen.

5.1 Notes for next year

Steven felt that this year had worked well and was an excellent template for from now on. Mark said we needed earlier advertising, and more ROs. Ben said we needed a person on the desk to handle entrys, and to get there earlier. A round of applause was

6 RO Dinner

given to John for getting Terry McAuley to present the prizes. Mark reported presenting Terry with a bottle of whiskey for having done this. Mark also pointed out that next year we need to have the perspex shield and a video camera to film the finals from in front of the firing line.

6 RO Dinner

The invitation list was discussed. Quick votes were taken. College dignitaries were not to be invited. Other shooting community people were to be invited, along with all ex-ROs and ex-Captains. John said Ray would have to help hunt down the ex-Captains. Steven suggested this was best saved for the 40th anniversary dinner.

John asked about the invitation list for the ex-Captains shoot held on the day after Boxing day. It was agreed that all ROs, ex-ROs, ex-Captains and Committee were to be invited.

Mark then reported that the booking for freemans on the 15th was impossible. The 18th was discussed but dropped due to large numbers of college functions on that day. The 17th was selected instead. Task: Book Freemans for the 17th (Mark)

7 Turkey Shoot

Ben has bought prizes and catering supplies for this event. Beer has been procured, but will have to be sold at £1 per can. Mugs/Cups are needed for the mulled wine. John has a set of 12 in his house awaiting transport to the range. It was suggested that Niamh bring them as she’s the stronger of the the pair. Task: Bring in mugs (John)

8 National Squad

The upcoming Dec 1st/2nd National Squad training session was discussed. As it was to be 10m Air Rifle only, few were interested. Peter, Mark and Ben were the only interested and qualified parties.

9 Christmas lists

Mark said he wanted to see DURC give christmas cards and presents to appropriate people. Superintendent Daniel Quinn, College Security, Terry McAuley, Trevor West and Pat Moray were to receive a bottle of whiskey, Cathy and Geraldine were to receive a box of chocolates or a bottle of wine, and charity christmas cards should be sent to all in DUCAC, the Department of Sport, and so forth. Ben will handle this as Mark is handling this for the NTSA. Task: Buy whiskey/wine/chocolate/cards for christmas lists (Ben)

10 Capital Grant

10 Capital Grant

Ben says we have 1600 in total, 300 for a gun safe and 1300 for a new air rifle. John says we are applying to the Department of Tourism and Sport’s Capital Grants program, so the money is frozen. DUCAC will be more flexible with funding if they see us as not fully dependant on them. It was noted that Terry McAuley and Pat Moray were not happy at the prospect of DURC acquiring more rifles.

11 Any Other Business

11.1 Jackets

Steven suggested getting Frank De Salle’s jacket converted from a Right-handed to an ambidexterous jacket at a cost of approx. £30. This was enthuastically agreed on and it was suggested we do this for all the other jackets, though it was felt doing it for a few was sufficent. John also suggested refitting all the jackets with new zips. Task: Follow-up on jacket conversions (Steven)

11.2 Pony Club coaching

Sunday is the pony club coaching day in westmeath. Peter and Mark will be at National Squad training. John is committed to spending the day with Niamh as the club has been getting all his weekend time of late. Conor is working. Deirdre was a bit hesitant as she felt underqualified. Ben will go, but not alone. Liam and Adrian were suggested, but Steven hadn’t asked them and wanted to get a definitive answer at the meeting. Peter agreed to go instead of going to squad training. Mark volunteered to go to the next coaching session, and gave dates when he was free. Task: Westmeath pony club coaching (Peter, Ben)

11.3 Student Newspapers

Steven said we were ready now to submit an article to the Student Record and the Trinity News. Mark said he thought we should submit every week. Steven thought that risked sounding boring after a while. Conor forwarded the complaints over the Rifle Club mailing list. Steven found this odd as he thought the list was running well with a low volume. Mark said that two complaints was not sufficent to shut down a service for 400. Alan suggested coordinating the postings to the list. Steven said this was difficult to arrange in reality. All post couldn’t be sent on one day. Mark pointed out that he’d had a half-dozen requests to change email addresses (indicating that the users were reading the posts and wished to continue to do so.).

11.4 Website

11 Any Other Business

Steven congradulated Mark and Conor on the website as it was now regularly updated with lots of useful information. The Photos section was the only problem. Alan said that from a home connection this page was unusable. Mark and Conor agreed, and there had been other complaints. Mark and Conor had a technical discussion on this problem, but the basic solution was not found. They will work on the problem. Task: Fix problems in photo pages on web (Conor,Mark)

11.5 AGM & Constitutional reform

Ben pointed out that the constitutional reform had not yet been dealt with. It was agreed a meeting of interested parties was needed, and soon so it wouldn’t be a lastminute affair. Steven suggested next week. Ben wanted the 40th anniversary discussion next week, but was overrulled.

11.6 Shooting nights

Deirdre wanted to have 3 .22 nights and 2 air nights a week. It was agreed we would change to this model when the air rifle season wound up in feburary.

11.7 Code of Conduct

Mark asked for a progress report on our Code of Conduct. No progress had been made. Steven took the NTSA Code of Conduct to use as a template for our Code which will be added to the club regulations. Task: Draft Code of Conduct (Steven)

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain

Mark Dennehy, Club Secretary

Tasks agreed on at meeting

Mark Look into the Regents House address (27 Nov 2001) [/138] . . . . . /2 Mark Recruit shooters for the postal competitions (27 Nov 2001) [/152] . /2 Alan Source and Price hardware for keyswitch (27 Nov 2001) [/266] … /2 Alan Announce RO Meeting (27 Nov 2001) [/295] . . . . . . . . . . . . . /2 John Adjust 603s trigger and report (27 Nov 2001) [/327] . . . . . . . . . /2 Mark Book Freemans for the 17th (27 Nov 2001) [/378] . . . . . . . . . . . /2 John Bring in mugs (27 Nov 2001) [/389] . . . . . . . . . . . . . . . . . . /2 Ben Buy whiskey/wine/chocolate/cards for christmas lists (27 Nov

2001) [/410] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Steven Follow-up on jacket conversions (27 Nov 2001) [/433] . . . . . . . . . /2 Peter, Ben Westmeath pony club coaching (27 Nov 2001) [/448] . . . . . . . . . /2 Conor,Mark Fix problems in photo pages on web (27 Nov 2001) [/476] . . . . . . /2 Steven Draft Code of Conduct (27 Nov 2001) [/499] . . . . . . . . . . . . . /2